OpenStudy (anonymous):

Write a radical in which the radicand consists of a coefficient and a variable. Use a coefficient that can be simplified best using the greatest perfect square method. Then, simplify the radical completely. You can use "sqrt()" to show a square root. For example, can be written as 7sqrt(a). Can someone help me with this? This is really confusing me, i'm not even sure I know what the question is asking exactly... I just don't understand it. Can someone please explain it to me?

jhonyy9 (jhonyy9):

hope that this will help you understanding - so radical indice 2 sign squareroot what is equal in exponentialy form by 1/2 - radical indice 3 is cuberoot what in exponent form is 1/3 - so for your help for example 4^(1/2 = (2^2)^(1/2) =2^(2*1/2) = =2^(2/2) =2^1 =2 - but this is squareroot of 4 =2 - ok ?

OpenStudy (anonymous):

is that the answer? the problem I worked out was like this: sqrt(8a) = sqrt(2*2^2*a) = 2sqrt(2a) but I'm not sure if that was correct, i'm literary going step by step from my notes... but is what you wrote the actual full answer to the problem? I hope you don't mind me being so specific, this unit is really tough for me... but thanks for your help too! :)

OpenStudy (anonymous):

wait, no the square root of 4, I just noticed that...

OpenStudy (anonymous):

sorry... lol

jhonyy9 (jhonyy9):

ok yw. was my pleasure - but for your exersice can you tel me how many will be cuberoot(8) = ?

jhonyy9 (jhonyy9):

or cuberoot64 = ?

jhonyy9 (jhonyy9):

so cuberoot8 = cuberoot(2^3) =(2^3)^(1/3 = 2^(3/3) = 2^1 =2 - can you do it this with cuberoot64 ?

OpenStudy (anonymous):

4? is that right? 4 x 4 = 16, and 16 x 4 = 64

OpenStudy (anonymous):

ohh wow, I don't think i'm getting this...

jhonyy9 (jhonyy9):

yes is right

OpenStudy (anonymous):

it is? wow I thought it was wrong... lol :)

jhonyy9 (jhonyy9):

5root32 = ?

OpenStudy (anonymous):

I got 28.284 on my calculator

jhonyy9 (jhonyy9):

why ? this sign radical indice 5 of 32

OpenStudy (anonymous):

well I just plugged in 5 squreroot of 32in my calculator and that was the answer, do you mean this way: (5/32)^1/5 0.689<-- this being the answer?

jhonyy9 (jhonyy9):

no check it please what i have wrote to you firstly in my answer radical indice 5 of 32 so this sign 32 on exponent 1/5 so what sign that you ned getting one number what multiplie 5 time will add to you 32

OpenStudy (anonymous):

nothing multiplied by 5 can get 32?

jhonyy9 (jhonyy9):

ok check you with 2

OpenStudy (anonymous):

ohhh wait,, wow I feel stupid, 2^5=32

OpenStudy (anonymous):

ugh, sorry... like I said, im not to good in this unit..

jhonyy9 (jhonyy9):

4root256 = ?

OpenStudy (anonymous):

64

OpenStudy (anonymous):

right?

jhonyy9 (jhonyy9):

why ?

OpenStudy (anonymous):

because, 4squareroot in 256= 64 on calc?

jhonyy9 (jhonyy9):

not is right how i have wrote before to you ?

jhonyy9 (jhonyy9):

not is right how i have wrote before to you ?

jhonyy9 (jhonyy9):

not is 4squreroot this is just 4root what sign radical indice 4

OpenStudy (anonymous):

its, 4! sorry, the way you wrote it made it seem like you were asking for 4squrerooted to 256, sorrry, like I said, much easier when its hand on and in person... but 4^4=256

OpenStudy (anonymous):

okay, do you think you could write out what we have done so far, maybe I can understand it better if I see the whole thing?? im still trying to organize it on paper... lol

jhonyy9 (jhonyy9):

ok after these can you tel me what is 9 on exponent 1/3

OpenStudy (anonymous):

9^1/3? is that what your asking? I got, 2.1for that... 2.0800 is exact from calc

OpenStudy (anonymous):

9/3 or 3?

OpenStudy (anonymous):

sorry for that first one, I got confused again...

OpenStudy (anonymous):

is that right?

jhonyy9 (jhonyy9):

no this sign cuberoot of 9

jhonyy9 (jhonyy9):

so can you calcule how many will be squarroot25

OpenStudy (anonymous):

umm i'm getting confused again, do you mean 9squareroot25? that would = 45?

jhonyy9 (jhonyy9):

ok 9sqrt25 yes is equal 45

OpenStudy (anonymous):

i'm so sorry for this, I guess I just don't get it... thats why I asked for you to write it out, I guess i'm thinking that if I see the whole problem together, I mite be able to see what i'm doing wrong...

OpenStudy (anonymous):

ohh that was what u were asking, wow sorry, again.. lol

jhonyy9 (jhonyy9):

do you understand it now sure right all ?

OpenStudy (anonymous):

I don't know if I do or not, I just follow the formulas and plug in the numbers, when it comes to stuff like this I get confused really easily...but i'm gonna give it my best shot, I'm pretty sure I understand what we have gone over, but it still may take me little so bare with me please...

OpenStudy (anonymous):

okay, the way I did this is, I just went back and wrote done exactly what we both discussed. so cuberoot8 = cuberoot(2^3) =(2^3)^(1/3 = 2^(3/3) = 2^1 =2 - can you do it this with cuberoot64 ? 4 x 4 = 16, and 16 x 4 = 64 5root32 = 2^5=32 4root256 = 4^4=256 what is 9 on exponent 1/3 sign cuberoot of 9 so can you calcule how many will be squarroot25 9sqrt25 = 45 im sorry to say that im still a little confused looking at this...

jhonyy9 (jhonyy9):

you wann calcule some exersices again ?

jhonyy9 (jhonyy9):

you wann calcule some exersices again ?

OpenStudy (anonymous):

I just want to get this problem over with, im sorry but this is really frustrating for me, i'm usually able to grasp algebraic concepts like these easily but I guess its just not my best subject... can we go back to the original question and work on that? I want to be able to understand the problem...

OpenStudy (anonymous):

you know, this one: Write a radical in which the radicand consists of a coefficient and a variable. Use a coefficient that can be simplified best using the greatest perfect square method. Then, simplify the radical completely. You can use "sqrt()" to show a square root. For example, can be written as 7sqrt(a). my example was this:sqrt(8a) = sqrt(2*2^2*a) = 2sqrt(2a) You wrote this: cuberoot8 = cuberoot(2^3) =(2^3)^(1/3 = 2^(3/3) = 2^1 =2 can you do it this with cuberoot64 ? and then I wrote: 4 x 4 = 16, and 16 x 4 = 64

OpenStudy (anonymous):

or should I do something different, maybe like this??: sqrt(125x^5) sqrt5 * sqrt25 * sqrtx^4 * sqrtx sqrt5 * 5 * x^2 * sqrtx 5x^2sqrt(5x)

jhonyy9 (jhonyy9):

yes is right cuberoot64 =cuberoot(4*4*4)=cuberoot(4^3)=4

OpenStudy (anonymous):

im just trying to make sure I understand the question, and that I don't sound like an idiot when I turn this in... lol :)

OpenStudy (anonymous):

okay so, would this be how you write that? sqrt(8a) = sqrt(2*2^2*a) = 2sqrt(2a) cube root8 = cuberoot(2^3) =(2^3)^(1/3 = 2^(3/3) = 2^1 =2 cuberoot64 =cuberoot(4*4*4)=cuberoot(4^3)=4

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

so that the answer? really?

OpenStudy (anonymous):

wow! thanks so much!

OpenStudy (anonymous):

Now that I can see the whole problem, it does make more sense, but i'm still gonna have to practice. You know, (i'm home schooled) I think that because I never have a teacher working with me in person(its all on the computer)that, that mite be why I have so much trouble with this, anyway thanks again, you were a really big help! :)

OpenStudy (anonymous):

okay, i'm gonna go submit that, thanks again and I mite be on later so if I need anymore help(and if your online) i'll let you know! :)

OpenStudy (anonymous):

I won't know what grade I'll get on the assignment until probably Monday, but again, I wanted to say thanks! you have no idea haw big of a help you were! :)

jhonyy9 (jhonyy9):

was my pleasure good luck bye